The next dividend payment by Hot Wings, Inc., will be $2.85 per share. The dividends are anticipated to maintain a 2 percent growth rate forever.

  

Required:
If the stock currently sells for $42 per share, what is the required return?

2.00%

8.35%

6.79%

8.61%

8.79%
Answer #1

Information provided:

Next dividend payment= $2.85

Growth rate= 2%

Current stock price= $42

The stock’s required return is calculated using the dividend discount model. It is calculated using the below formula:

Ke=D1/Po+g

Where:

D1= Next year’s dividend

Po=Current stock price

g=Firm’s growth rate

Ke= $2.85/ $42 + 0.02

     = 0.0679 + 0.02

     = 0.0879*100

     = 8.79%.

Hence, the answer is option e.

In case of any query, kindly comment on the solution.